OpenStudy (anonymous):

2. Solve for g. √6g - 23 = √12 - g 5 7 11 15

OpenStudy (anonymous):

5

OpenStudy (anonymous):

thanks! can you help with another question?

OpenStudy (anonymous):

yea, show

OpenStudy (anonymous):

solve for d. √d^2-11=5

OpenStudy (anonymous):

good question i think it will be five

OpenStudy (anonymous):

sorry i forgot to put the choices. 4, -4 5, -5 5, 6 6, -6

OpenStudy (anonymous):

\[\sqrt{d^2-11}=5 ; d^2-11 = 25 ; d = +-6\]

OpenStudy (anonymous):

d

OpenStudy (anonymous):

thanks so much! can you help with one more please??

OpenStudy (anonymous):

ok, next ?

OpenStudy (anonymous):

which is the extraneous solution of -x = √2x+15?

OpenStudy (anonymous):

i don't understand ? i know a little english

OpenStudy (anonymous):

extraneous solution ?

OpenStudy (anonymous):

an extraneous solution is an apparent solution that does not satisfy the original equation.

OpenStudy (anonymous):

solve equation?

OpenStudy (anonymous):

the choices are -5 -3 5 3

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

another question ?

OpenStudy (anonymous):

what is the domain of the function y = 3 √6x+42? x≥0 x≤7 x≥-6 x≥-7

OpenStudy (anonymous):

d

OpenStudy (anonymous):

next ?

OpenStudy (anonymous):

thanks!